Kimberly took three hours to drive to her friends home on the weekend. Her average speed was 85 km/h for the first two hours. What was her average speed for the last hour of the trip if her average speed was 75 km/h for the whole trip?
Answer by ankor@dixie-net.com(22740)

Kimberly took three hours to drive to her friends home on the weekend.
Her average speed was 85 km/h for the first two hours.
What was her average speed for the last hour of the trip if her average speed was 75 km/h for the whole trip?
:
Let s = the speed on the last hour of the trip
:
Write a distance equation, dist = time * speed
:
dist 1st 2 hr + dist last hr = total dist
2(85) + 1s = 3(75)
190 + s = 225
s = 225 - 190
s = 35 km/h for the last hour